Hillside issues $1,700,000 of 8%, 15-year bonds dated January 1, 2017, that pay interest semiannually on June 30 and December 31 The bonds are issued at a price of $2,080,794 Required 1. Prepare the January 1, 2017, journal entry to record the bonds issuance 2(a) For each semiannual period, complete the table below to calculate the cash payment. 2(b) For each semiannual period, complete the table below to calculate the straight-line premium amortization. 2(c) For each semiannual period, complete the table below to calculate the bond interest expense 3. Complete the below table to calculate the total bond interest expense to be recognized over the bonds life 4. Prepare the first two years of an amortization table using the straight-line method 5. Prepare the journal entries to record the first two interest payments. Answer is not complete Complete this question by entering your answers in the tabs below Req 1 Req 2A to 2C Req 3 Req 4 Req 5 Complete the below table to calculate the total bond interest expense to be recognized over the bonds life Total bond interest expense over life of bonds Amount repaid: 30payments of Par value at maturity Total repaid Less amount borrowed Total bond interest expense 1,700,000 1,700,000 2,080,794 $ (380,794) Req 2A to 2C Req 4

Req 1 Req 2A to 2C Req 3 Req 4 Req 5 Prepare the first two years of an amortization table using the straight-line method Semiannual Period-End 01/01/2017 06/30/2017 12/31/2017 06/30/2018 12/31/2018 Unamortized arrying Value Premium $380,794 $ 2,080,794 350,330 368,101 355,408 2,068,101 2,055,408 2,042,715 2,030,021 Req 3 Req 5 >

Journal entry worksheet 2 Record the first interest payment on June 30, 2017 Note: Enter debits before credits. Date General Journal Debit Credit Jun 30, 2017 Bond interest expense Cash

Journal entry worksheet Record the second interest payment on December 31, 2017 Note: Enter debits before credits. Date General Journal Debit Credit Dec 31, 2017 68,000 Record entry Clear entry View general journal

Please help me with the remaining questions I have left
Answer #1

Par maturity value =1700000

Annual value =8%

Semi annual cash payment (1700000*8%*6/12)=68000

Total bond interest expense over life of bonds
Amount repaid
30 payments of 68,000 2,040,000
par value at maturity 1,700,000
Total repaid 3,740,000
Less amount borrowed 2,080,794
Total bond interest expense. 1,659,206

Semiannual Period-EndUnamortized premiumCarrying Value 1/1/2017 06/30/2017 12/31/2017 06/30/2018 12/31/2018 $380,794 368,101 355,408 342,715 330,022 $2,080,794 2,068,101 2,055,408 2,042,715 2,030,022Par (maturity) value Bonds price Premium on Bonds Payable(1700000-2080794) 1,700,000 2,080,794 380794 12693 Straight line preUnamortized premium on 6/30/2017 =380794-12693=368101,similar calculation is applied for remaining columns ie, unamortized premium less straight line premium amortization. In the case of carrying value also same logic is used.

Date General Journal Debit Credit
Jun 30, 2017 Bond interest expense 55,307
Premium on bonds payable 12693
Cash 68000
Dec 31, 2017 Bond interest expense 55307
Discount on bonds payable 12693
Cash 68000

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW. PLEASE RATE MY ANSWER BY HITTING ??